# On the Generalization of SFT: A Reinforcement Learning Perspective with Reward Rectification

### v2

@@ 4.3 Exploratory Experiment - Code Generation Task
To further validate the effectiveness of DFT, we investigate its performance on another reasoning-intensive task: code generation.
4.3.1 Setup and Implementation Details
We adopt UltraFeedback (Cui et al., 2024) as the training dataset. From this corpus, we sample 10,000 prompts and, for each prompt, select the response with the highest average score to perform supervised fine-tuning (SFT) (Du et al., 2025). Model performance is assessed on three widely used code generation benchmarks: HumanEval (Chen et al., 2021), HumanEval+ (Liu et al., 2023), and MultiPL-E (Cassano et al., 2023). Training is conducted for one epoch with a learning rate of $5\times 10^{-5}$, a warm-up ratio of 0.05, and a batch size of 16.
4.3.2 Results
@@
HumanEval HumanEval+ Python C++ Java PHP TS C# Bash JS Avg.
@@
4.4 Exploratory Experiment - Multi-Modal Reasoning
To further investigate the generalization of DFT, we extend our evaluation to multi-modal reasoning tasks that require both visual understanding and mathematical problem solving.
@@ 4.4 Exploratory Experiment - Multi-Modal Reasoning
4.4.1 Setup and Implementation Details
We use the WeThink dataset (Yang et al., 2025) for training. The model is fine-tuned using LLaMA-Factory (Zheng et al., 2024) and evaluated with VLMEvalKit (Duan et al., 2024). We train the model for 1 epoch with a learning rate of 5e-5. To comprehensively assess reasoning capabilities, we adopt a suite of multi-modal reasoning benchmarks including MathVerse (Zhang et al., 2024a), MathVision (Wang et al., 2024), and WeMath (Qiao et al., 2024) for evaluation.
4.4.2 Results
@@
4.5 Analysis
@@ 5 Conclusion
In this work, we address the well-documented generalization gap between Supervised Fine-Tuning and Reinforcement Learning. We provide a novel theoretical analysis, demonstrating that the standard SFT gradient is equivalent to a policy gradient update with an ill-posed implicit reward that is inversely proportional to the model’s confidence. This insight explains SFT’s tendency to overfit and its unstable optimization dynamics. Building on this analysis, we introduce DFT, a simple yet powerful method that rectifies this issue by dynamically reweighting the SFT loss with the token probability. This one-line modification stabilizes the learning process and promotes better generalization. Our extensive experiments show that DFT consistently outperforms standard SFT across various models and challenging mathematical reasoning benchmarks. Furthermore, when adapted to an offline RL setting, DFT  surpasses established online and offline RL algorithms, highlighting its effectiveness and efficiency. Our work provides both a deeper understanding of SFT and a practical, high-impact solution that significantly closes the performance gap with more complex RL methods.
@@ A.1 Usage of LLM
A.2 Limitations
While our experiments demonstrate the effectiveness of DFT on mathematical reasoning benchmarks and code generation tasks, the evaluation scope remains limited. We have not yet assessed its performance on broader task categories or with larger-scale LLM, which we leave for future exploration. Moreover, DFT may not offer universal benefits across all scenarios. In domains that primarily involve the acquisition of factual knowledge, conventional SFT may remain the most efficient approach. Our aim is not to assert that DFT universally outperforms SFT, but rather to offer a new perspective on objective design by analyzing the distinction between RL and SFT.
A.3 Detailed Derivation of Equation (5)
@@ A.4 Discussions and Insights
DFT offers a simple yet effective approach, prompting us to reflect on why it might actually work. One intuitive explanation lies in its ability to learn from noisy data (Freund, 2009). Sasaki & Yamashina (2020) propose an imitation learning algorithm for learning from noisy demonstrations, based on the core idea of avoiding the fitting of data that is difficult to model, as such data is likely to originate from suboptimal behaviors, i.e., noise. Their method introduces a weighted behavioral cloning objective, where the weights are derived from a previously trained policy’s confidence in each action. Similarly, the weighting mechanism in DFT  shares the same intuition, but instead of relying on a fixed old policy model to compute confidence scores, it uses a single policy model to perform confidence-based weighting on-the-fly during training.
A.5 Comparision with Concurrent work iw-SFT
We include a concurrent method, Importance-Weighted SFT (iw-SFT) (Qin & Springenberg, 2025), for comparison. All training settings follow those reported in the original paper, except that we set the number of training epochs to 1.
@@ A.5 Comparision with Concurrent work iw-SFT
A.6 Exploratory Experiment - OpenR1-Math Training Dataset
@@
A.7 Exploratory Experiment - PEFT Training Setting
@@
A.8 Training Hyper-Parameters Ablation

### v3

@@ 4.3 Exploratory Experiment - Code Generation Task
HE HE+ Python C++ Java PHP TS C# Bash JS Avg.
@@
Implementation details.
We adopt UltraFeedback (Cui et al., 2024) as the training dataset. From this corpus, we sample 10,000 prompts and, for each prompt, select the response with the highest average score to perform supervised fine-tuning (SFT) (Du and others, 2025). Model performance is assessed on three widely used code generation benchmarks: HumanEval (Chen et al., 2021), HumanEval+ (Liu et al., 2023), and MultiPL-E (Cassano et al., 2023). Training is conducted for one epoch with a learning rate of $5\times 10^{-5}$, a warm-up ratio of 0.05, and a batch size of 16.
4.4 Exploratory Experiment - Multi-Modal Reasoning
Implementation details.
We use the WeThink dataset (Yang et al., 2025) for training. The model is fine-tuned using LLaMA-Factory (Zheng et al., 2024) and evaluated with VLMEvalKit (Duan et al., 2024). We train the model for 1 epoch with a learning rate of 5e-5. To comprehensively assess reasoning capabilities, we adopt a suite of multi-modal reasoning benchmarks including MathVerse (Zhang et al., 2024b), MathVision (Wang et al., 2024), and WeMath (Qiao et al., 2024) for evaluation.
@@ 4.4 Exploratory Experiment - Multi-Modal Reasoning
4.5 Limitations of DFT: A Case Study on Factual Knowledge
While DFT consistently outperforms SFT on reasoning-heavy tasks, it may not always be the better choice, particularly in factual knowledge domains. We conduct an exploratory experiment on the Natural Questions dataset (Kwiatkowski et al., 2019), which consists of real-user, open-domain factual queries grounded in Wikipedia articles.
In this setting, we find that SFT improves performance from 31.24% to 36.62%, while DFT unexpectedly reduces it to 30.14%. This result reveals an important limitation of DFT: because it reweights samples based on the model’s own confidence, it tends to reinforce the model’s existing beliefs. When the model lacks sufficient factual knowledge, such reinforcement may hinder effective learning instead of facilitating it.
This case suggests that DFT is most effective when the task aligns well with the model’s prior competence, such as logical reasoning or structured prediction. In contrast, when the objective is to absorb new factual information, especially in domains beyond the model’s current capabilities, SFT remains a more reliable and stable fine-tuning strategy.
Table 5: Comparison of weighting strategies on mathematical reasoning benchmarks.
Math500 Minerva Math Olympiad Bench AIME24 AMC23 Avg.
Qwen2.5-Math-1.5B 31.66 8.51 15.88 4.16 19.38 15.92
Sentence-Level Weighting 31.26 8.05 16.47 3.12 19.84 15.75
Geometric-Mean Weighting 42.87 12.34 13.03 1.23 16.56 17.21
Token-Level Weighting 64.89 20.94 27.08 6.87 38.13 31.58
4.6 An Empirical Comparison with Sentence-Level Weighting
Our framework applies confidence-based weighting at the token level. While this design was primarily motivated by numerical stability, we also compared it against two sentence-level variants to better understand their behavior.
The first variant uses the full sequence probability to scale the loss. However, these values are extremely small in practice, making the loss nearly uninformative and producing a highly skewed weight distribution that is difficult to tune. To address this, we also evaluated a geometric-mean variant inspired by GSPO (Zheng et al., 2025), which rescales sentence probabilities to avoid numerical collapse. Although this version is more stable, it still provides a weak training signal and offers limited performance gains.
As shown in Table 5, both sentence-level strategies lead to minimal changes over the base model, while our token-level formulation delivers substantial and consistent improvements, raising average accuracy from 15.92 to 31.58. These results demonstrate that token-level weighting provides a more reliable optimization signal and significantly stronger empirical performance.
4.7 Analysis of Probabilities
@@ 5 Conclusion
In this work, we revisit the well-known generalization gap between SFT and RL. We offer a theoretical perspective showing that the standard SFT gradient can be interpreted as a policy gradient with an ill-posed, implicitly defined reward inversely related to model confidence. This formulation helps explain the instability and limited generalization observed in SFT training. Motivated by this analysis, we introduce DFT, a simple yet effective method that dynamically reweights the SFT loss using the token probability. This one-line change improves gradient stability and leads to better generalization. Our empirical results show that DFT consistently improves over standard SFT across a range of models and challenging mathematical reasoning tasks. Beyond supervised settings, we adapt DFT to offline RL scenarios and find that it outperforms several established online and offline RL baselines, suggesting broader applicability. Overall, this work contributes both a refined understanding of SFT’s limitations and a lightweight, practical method that helps bridge the gap to more complex RL-based approaches.
Limitations.
While our experiments demonstrate the effectiveness of DFT on mathematical reasoning benchmarks and code generation tasks, the evaluation scope remains limited. We have not yet assessed its performance on broader task categories or with larger-scale LLM, which we leave for future exploration. Moreover, DFT can not offer universal benefits across all scenarios. In domains that primarily involve the acquisition of factual knowledge, conventional SFT still remains the most efficient approach. DFT may also not be an ideal choice for hard examples or domains underrepresented in the training data, since it assigns low initial probabilities to such samples, reducing their learning weight. Our aim is not to assert that DFT universally outperforms SFT, but rather to offer a new perspective on objective design by analyzing the distinction between RL and SFT. Besides, an important future direction is to explore non-uniform or quality-aware reward assignments for demonstrations.
Acknowledgements
Supported by Jiangsu Province Carbon Peak Carbon Neutrality Science and Technology Innovation Special Fund Project (Grant No. BT2025029), National Natural Science Foundation of China (Grant No. 62576091), and Big Data Computing Center of Southeast University.
Ethics Statement
This work adheres to the ICLR Code of Ethics. Our study does not involve human subjects, personally identifiable information, or proprietary data. All datasets used, including NuminaMath, OpenR1-Math, UltraFeedback, and WeThink, are publicly available and documented in the appendix. The proposed method is a simple training strategy that modifies gradient computation for improved generalization. It does not introduce any new capabilities that could cause harm, nor does it enable misuse beyond the standard capabilities of existing large language models. We are not aware of any potential risks related to bias, fairness, or security that arise specifically from the method proposed. Nonetheless, we acknowledge that like any fine-tuning strategy, DFT may inherit biases present in the underlying data or model, and future research may explore safeguards for these scenarios. No conflicts of interest, legal compliance issues, or sponsorship-related influences are present in this work.
Reproducibility Statement
We have taken multiple steps to ensure the reproducibility of our work. All datasets used in our experiments are publicly available and properly cited in the main text and appendix. Training configurations, including model architectures, hyperparameters, optimizers, and evaluation settings, are described in detail in Section 4 and Appendi A.5-A.6. Theoretical claims, including the equivalence between SFT and policy gradient, are formally derived in Appendix A.2. Experimental results include multiple model scales, tasks, and training settings to validate robustness. A complete implementation of our method is included in the supplementary material, along with scripts for reproducing all reported results. We will release the full source code and training logs upon publication to further support reproducibility.
@@ A.1 Usage of LLM
A.2 Detailed Derivation of Equation (5)
@@ A.3 Discussions and Insights
DFT offers a simple yet effective approach, prompting us to reflect on why it might actually work. One intuitive explanation lies in its ability to learn from noisy data (Freund, 2009). Sasaki and Yamashina (2020) propose an imitation learning algorithm for learning from noisy demonstrations, based on the core idea of avoiding the fitting of data that is difficult to model, as such data is likely to originate from suboptimal behaviors, i.e., noise. Their method introduces a weighted behavioral cloning objective, where the weights are derived from a previously trained policy’s confidence in each action. Similarly, the weighting mechanism in DFT  shares the same intuition, but instead of relying on a fixed old policy model to compute confidence scores, it uses a single policy model to perform confidence-based weighting on-the-fly during training.
A.4 Comparision with Concurrent work iw-SFT
We include a concurrent method, Importance-Weighted SFT (iw-SFT) (Qin and Springenberg, 2025), for comparison. All training settings follow those reported in the original paper, except that we set the number of training epochs to 1.
@@ A.4 Comparision with Concurrent work iw-SFT
A.5 Exploratory Experiment - OpenR1-Math Training Dataset
@@
A.6 Exploratory Experiment - PEFT Training Setting
@@
A.7 Training Hyper-Parameters Ablation